40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 873     Entry time: Wed Mar 7 11:39:37 2012
Author: Tara, Frank 
Type: Notes 
Category: TempCtrl 
Subject: acav temp tuning - 2nd try 

previous measurement is invalid - due to changing in the alignment from floated table to unfloated table over night we don't know what we really measured.

What we can say is that the beat frequency changed from 181.265MHz to 185.9MHz, so by 4.66MHz for a 0.2K increase of the shield using 1.53V heater voltage.

aligned everything properly, locked both cavities, changed tuning input range to 100kHz and turned heater off. We might stop the new measurement at any time as we got the basic information we need

ELOG V3.1.3-